OpenStudy (anonymous):

Transform the following polar equation into an equation in rectangular coordinates: r=2 cos theta x = 2 (x-1)^2+y^2=1 y = 2x x + y = 2

OpenStudy (anonymous):

r=2 cos theta multiply r to both sides: r^2 = 2 rcos(theta) since r^2 = x^2 + y^2 and x = rcos(theta), x^2 + y^2 = 2x can you take it from here?

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I thought i understood it but i still need the answer to make sure

OpenStudy (eyust707):

in polar x= rcostheta and y = rsintheta

OpenStudy (eyust707):

you can just sub those in and your good to go

OpenStudy (anonymous):

are you able to get here: x^2 + y^2 = 2x ?